OpenStudy (anonymous):

WILL MEDAL AND FAN!!!!!!!!!!!!!! 6x^2-15=x

OpenStudy (faiqraees):

Take x on left hand side and then use the quadratic formula to work the factors

OpenStudy (anonymous):

@ethane

OpenStudy (ethane):

lol you want my help?

OpenStudy (anonymous):

Yeah lol. I did what he said and still didn't get the right answer.

OpenStudy (faiqraees):

Can you show your working?

OpenStudy (ethane):

hmm...ok ill help you

OpenStudy (anonymous):

Thank ya

OpenStudy (ethane):

ok so you need to get the xs to one side

OpenStudy (ethane):

crap this is hard XD

OpenStudy (ethane):

@smile.89745

OpenStudy (faiqraees):

OKay so first start by taking the whole equation on one side 6x^2-x-15 = 0 Now this is in the form on a quadratic equation and hence the quadratic formula can be applied \[\large\rm x=\frac{-b±\sqrt{b^2-4ac}}{2a} \text{where a, b and c denotes the coefficients} \]\[\large\rm \text{ of ax²+bx+c=0} \]

OpenStudy (scoobydo):

this will help Simplifying 6x2 + -15 = x Reorder the terms: -15 + 6x2 = x Solving -15 + 6x2 = x Solving for variable 'x'. Reorder the terms: -15 + -1x + 6x2 = x + -1x Combine like terms: x + -1x = 0 -15 + -1x + 6x2 = 0 Factor a trinomial. (-3 + -2x)(5 + -3x) = 0 Subproblem 1 Set the factor '(-3 + -2x)' equal to zero and attempt to solve: Simplifying -3 + -2x = 0 Solving -3 + -2x = 0 Move all terms containing x to the left, all other terms to the right. Add '3' to each side of the equation. -3 + 3 + -2x = 0 + 3 Combine like terms: -3 + 3 = 0 0 + -2x = 0 + 3 -2x = 0 + 3 Combine like terms: 0 + 3 = 3 -2x = 3 Divide each side by '-2'. x = -1.5 Simplifying x = -1.5 Subproblem 2 Set the factor '(5 + -3x)' equal to zero and attempt to solve: Simplifying 5 + -3x = 0 Solving 5 + -3x = 0 Move all terms containing x to the left, all other terms to the right. Add '-5' to each side of the equation. 5 + -5 + -3x = 0 + -5 Combine like terms: 5 + -5 = 0 0 + -3x = 0 + -5 -3x = 0 + -5 Combine like terms: 0 + -5 = -5 -3x = -5 Divide each side by '-3'. x = 1.666666667 Simplifying x = 1.666666667Solution x = {-1.5, 1.666666667}

OpenStudy (scoobydo):

did it help

OpenStudy (anonymous):

Yes thank you

OpenStudy (smile.89745):

yea idk @ethane

OpenStudy (smile.89745):

im not a math genius @ethane